What is the pH solution containing 0.30 M HCOOH and 0.52 M HCOOK?Compare your result with the pH of a 0.30 M HCOOH solution


1
Expert's answer
2022-03-10T10:47:43-0500

HCOOH(aq) + H2O(l) = H3O+(aq) + HCOO-(aq)

pKa = 3.77

"pH = pKa + log ( \frac{[HCOO^-]}{[HCOOH]} )"

"pH = 3.77 + log ( \frac{[0.52]}{[0.30]} )"

pH = 4.01


"Ka=\frac{[HCOO^-][H^+]}{[HCOOH]}=\frac{x^2}{0.3}"

"pKa=logx^2-log0.3"

"logx^2=0.52+3.77=4.26"

logx=pH=4.26/2=2.13